OpenStudy (anonymous):

solve by the elimination process 2x + 3y = 4 4x + 6y = 8 what is the solution of the system

OpenStudy (anonymous):

solve by the elimination process 2x + 3y = 4 4x + 6y = 8 what is the solution of the system

jhonyy9 (jhonyy9):

x=-1 y=2

OpenStudy (anonymous):

hi jhonyy9 can u show me how u arrived at this answer please or someone

OpenStudy (anonymous):

the reason i'm asking is because this is what i came up with: 2x + 3y = 4 -- if you want to eiminate x here, multiply by 2 4x + 6y = 8 (2x + 3y = 4)2 4x + 6y = 8 4x + 6y = 8 4x + 6y = 8 0 + 0 + 0 from this there is no solution unless I did something wrong

jhonyy9 (jhonyy9):

2x+3y=4 --- 2(-1)+3(2)=4 --- (-2)+6=4 --- 4=4 4x+6y=8 --- 4(-1)+6(2)=8 --- (-4)+12=8 --- 8=8